OpenStudy (anonymous):

The Charges q1= -18.5 x10-6 C and q2= 24.0 x 10-6 C are fixed at point O and P respectively on the base corners of a trapezoid. a) Compute the potential at points M and N. b) What is the potential difference between points N and M. c) Calculate the work done to move a point charge of 2.5x10-6 C at constant speed from point M to point N,

OpenStudy (anonymous):



OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@shifuyanli @Hotchellerae21

OpenStudy (michele_laino):

the electric potential at point M, is given by the subsequent formula: \[\Large V\left( M \right) = K\frac{{{q_1}}}{{0.2}} + K\frac{{{q_2}}}{d}\]

OpenStudy (anonymous):

what should be the label on the drawing?

OpenStudy (michele_laino):

where d is: